News & Annoucements

The situation at LSC

12 May 2017

We have had to discontinue observations with the telescope in Dome C at our Chilean site (LSC). The Sinistro camera in that dome (fl04) appeared to have a steep drop in temperature on 2017-05-07. The drop was actually a symptom of a problem with a circuit boards on the camera controller. A replacement board has been sent to Chile; the camera will be offline until the swap can be made.
The noise problem with the camera in Dome A (fl15) is still being investigated. (I reported this on April 26.) Consequently, we are currently only able to make observations with one telescope at our Chilean site: the fl03 Sinistro camera in Dome B. 
We recently re-processed all fl03 data that were acquired between 2017-03-21 and 2017-05-08. In the initial processing of these data, we accidentally introduced a slight (few %) error in the gains that were applied to quadrants 2,3,4. We detected the problem about two weeks ago, but it took time for us to devise and deploy the remedy. If you acquired images with the fl03 camera during this period, I encourage you to retrieve the corrected images from the archive.
A reminder: a link to our Network Operational Status page is in the footer of most of the pages on the LCO website. We do our best to keep the page up-to-date.

Return to News